3x/ (x^2+5x+6) = 5x/ (x^2+2x-3) - 2/(x^2+x-2)
Answer by lynnlo(4176)

3x(x-1)=5x(x+2)-2(x+3)
3x(x-1)-5x(x+2)+2(x+3)=0
-2x^2-11x+6=0
-((x+6)(2x-1))=0
(x+6)(2x-1)=0
x+6=0 OR 2x-1=0
x=-6 OR 2x-1=0
x=-6 OR 2x=1
DIVIDE BOTH SIDES BY 2
x=-6 OR x=1/2